Use the data above to calculate the amount of energy released (in kilojoules per milliliter) from the combustion of cooking fat:
Answer #1
C51H88O6 + 70 O2 = 51 CO2 + 44 H2O
Consider 1 mole of C51H88O6
Mass = moles x molar mass
= 1 x 797.246 = 797.246 g
Volume = mass/density
= 797.246/0.94 = 848.134 mL
Amount of energy released = -?H/volume of 1 mole
= 1310/848.134 = 1.54 kJ/mL
answered by: juan-luca